Tarnishing of silver is caused by small amounts of hydrogen sulfide released into the atmosphere by decaying organic material and volcanic activity. According to the following equation:


 4Ag(s) + 2H2S(g) + O2(g) → 2Ag2S(s) + 2H2(l)


How many grams of Ag2S can be produced from 1.70g H2S if sufficient amount of the other reactants are present?



1
Expert's answer
2021-01-15T06:38:59-0500

The molar masses are:


"M_{Ag_2S}=247.8g\/mol"

"M_{H_2S}=34.1g\/mol"


Therefore,

"m_{Ag_2S}=1.70g_{H_2S}\times\frac{1mol_{H_2S}}{34.1g_{H_2S}}\times\frac{2mol_{Ag_2S}}{2mol_{H_2S}}\times\frac{247.8g_{Ag_2S}}{1mol_{Ag_2S}}=12.4g"

Answer: 12.4 g